Operations Report

Crew 222 Operations Report 02/24/2020

SOL: 8

Name of person filing report: Valentin PRUDHOMME

Non-nominal systems: N/A

Notes on non-nominal systems: N/A

Sunday generator:
Off at 6:30am
On at 9:00pm SOC was 83%

Monday generator:
Off at 6:00am
On at 9:00 pm SOC was XX%

Diesel Reading – 74%

Station Propane Reading – 29% (refill asked)

Ethanol Free Gasoline – 0.5 gallons

Water (loft tank): 48 gallons

Water Meter: 149453.58 units

Water (static tank) – 304 gallons

Static to Loft Pump used – yes

Water in GreenHab – 72 gallons

Water in ScienceDome: 0 gallons

Toilet tank emptied: no

Deimos rover used: no
Hours: 146.2
Beginning charge: N/A
Ending charge: N/A
Currently charging: yes

Sojourner rover used: ASSIGNED TO DIRECTOR
Hours: N/A
Beginning charge: N/A
Ending charge: N/A
Currently charging: N/A

Spirit rover: no
Hours: 132.4
Beginning charge: N/A
Ending charge: N/A
Currently charging: yes

Opportunity rover used: no
Hours: 69.9
Beginning charge: N/A
Ending charge: N/A
Currently charging: yes

Curiosity rover used: yes
Hours: 138.5
Beginning charge: 100%
Ending charge: 94%
Currently charging: yes

Notes on rovers: N/A

ATV’s Used: yes
Reason for use: Run during the engineer EVA for the battery.
Oil Added? no
ATV Fuel Used: N/A
Hours the ATVs were Used today: 10 minutes
Notes on ATVs: The second one from the Hab doesn’t have anymore battery.
The red one was very difficult to start but I manage to do it.

HabCar used and why, where? The habcar is now in the shop awaiting repairs.

CrewCar used and why, where? N/A

General notes and comments: N/A

Summary of internet: All ok.

Summary of suits and radios: Suits 8 and 9 used and radio tested for the
EVA. all ok.

Summary of Hab operations: N/A

Summary of GreenHab operations: N/A

Summary of ScienceDome operations: N/A

Summary of RAM operations: After the rain of yesturday, the RAM wasn’t
waterproof and leaked on our electronics.

Summary of any observatory issues: N/A

Summary of health and safety issues: N/A

Questions, concerns and requests to Mission Support: N/A

Operations Report

Crew 222 Operations Report 02/23/2020

SOL: 7

Name of person filing report: Valentin PRUDHOMME

Non-nominal systems: N/A

Notes on non-nominal systems: N/A

Saturday generator:
Off at 5:30am
On at 5:00pm SOC was 80%

Sunday generator:
Off at 6:30am
On at 9:00pm SOC was XX%

Diesel Reading – 74%

Station Propane Reading – 30% (refill asked)

Ethanol Free Gasoline – 0.5 gallons

Water (loft tank): 47 gallons

Water Meter: 149426.79 units

Water (static tank) – 319 gallons

Static to Loft Pump used – yes

Water in GreenHab – 80 gallons

Water in ScienceDome: 0 gallons

Toilet tank emptied: no

Deimos rover used: no
Hours: 146.2
Beginning charge: N/A
Ending charge: N/A
Currently charging: yes

Sojourner rover used: ASSIGNED TO DIRECTOR
Hours: N/A
Beginning charge: N/A
Ending charge: N/A
Currently charging: N/A

Spirit rover: no
Hours: 132.4
Beginning charge: N/A
Ending charge: N/A
Currently charging: yes

Opportunity rover used: no
Hours: 69.9
Beginning charge: N/A
Ending charge: N/A
Currently charging: yes

Curiosity rover used: yes
Hours: 138.4
Beginning charge: 100%
Ending charge: 90%
Currently charging: yes

Notes on rovers: N/A

ATV’s Used: yes
Reason for use: Run during the engineer EVA for the battery.
Oil Added? no
ATV Fuel Used: N/A
Hours the ATVs were Used today: 10 minutes
Notes on ATVs: The second one from the Hab doesn’t have anymore battery.

HabCar used and why, where? The habcar broke down on 02/18/2020. It’s
parked at Bull Mountain Market.

CrewCar used and why, where? N/A

General notes and comments: N/A

Summary of internet: All was fixed yesterday evening. All ok now.

Summary of suits and radios: Suits 8 and 9 used and radio tested for the
EVA. all ok.

Summary of Hab operations: N/A

Summary of GreenHab operations: N/A

Summary of ScienceDome operations: N/A

Summary of RAM operations: N/A

Summary of any observatory issues: N/A

Summary of health and safety issues: N/A

Questions, concerns and requests to Mission Support: N/A

Operations Report

Crew 222 Operations Report 02/22/2020

SOL: 6

Name of person filing report: Valentin PRUDHOMME

Non-nominal systems: N/A

Notes on non-nominal systems: N/A

Friday generator:
Off at 6:30am
On at 9:00pm SOC was 80%

Saturday generator:
Off at 5:30am
On at 5:00pm SOC was 81%

Diesel Reading – 74%

Station Propane Reading – 31% (refill asked)

Ethanol Free Gasoline – 0.5 gallons

Water (loft tank): 50 gallons

Water Meter: 149408.46 units

Water (static tank) – 339 gallons

Static to Loft Pump used – yes

Water in GreenHab – 89 gallons

Water in ScienceDome: 0 gallons

Toilet tank emptied: no

Deimos rover used: no
Hours: 146.2
Beginning charge: N/A
Ending charge: N/A
Currently charging: yes

Sojourner rover used: ASSIGNED TO DIRECTOR
Hours: N/A
Beginning charge: N/A
Ending charge: N/A
Currently charging: N/A

Spirit rover: no
Hours: 132.4
Beginning charge: N/A
Ending charge: N/A
Currently charging: yes

Opportunity rover used: no
Hours: 69.9
Beginning charge: N/A
Ending charge: N/A
Currently charging: yes

Curiosity rover used: yes
Hours: 138.3
Beginning charge: 100%
Ending charge: 94%
Currently charging: yes

Notes on rovers: N/A

ATV’s Used: yes
Reason for use: Run during the engineer EVA for the battery.
Oil Added? no
ATV Fuel Used: N/A
Hours the ATVs were Used today: 5 minutes
Notes on ATVs: The second one from the Hab doesn’t have anymore battery.
It’s planned to be fixed by Dr Rupert on Sunday.

HabCar used and why, where? The habcar broke down on 02/18/2020. It’s
parked at Bull Mountain Market.

CrewCar used and why, where? N/A

General notes and comments: N/A

Summary of internet: Since 10am this morning, we can connect to the wifi
but it directly opens an internet page saying : "you failed to access to
the web page that you want to view." no matter the page we want to load.

Summary of suits and radios: Suits 7 and 9 used and radio tested for the
EVA.

Summary of Hab operations: N/A

Summary of GreenHab operations: N/A

Summary of ScienceDome operations: N/A

Summary of RAM operations: N/A

Summary of any observatory issues: N/A

Summary of health and safety issues: N/A

Questions, concerns and requests to Mission Support: N/A

Operations Report – February 25

Crew 222 Operations Report 02/25/2020

SOL: 9

Name of person filing report: Valentin PRUDHOMME

Non-nominal systems: N/A

Notes on non-nominal systems: N/A

Monday generator:
Off at 6:00am
On at 9:00pm SOC was 85%

Tuesday generator:
Off at 5:30am
On at 8:30pm SOC was XX%

Diesel Reading – 72%

Station Propane Reading – 27% (refill asked)

Ethanol Free Gasoline – 0.5 gallons

Water (loft tank): 47 gallons

Water Meter: 149478.07 units

Water (static tank) – 280 gallons

Static to Loft Pump used – yes

Water in GreenHab – 58.5 gallons

Water in ScienceDome: 0 gallons

Toilet tank emptied: yes

Deimos rover used: no
Hours: 146.2
Beginning charge: N/A
Ending charge: N/A
Currently charging: yes

Sojourner rover used: ASSIGNED TO DIRECTOR
Hours: N/A
Beginning charge: N/A
Ending charge: N/A
Currently charging: N/A

Spirit rover: yes
Hours: 132.8
Beginning charge: 100%
Ending charge: 58%
Currently charging: yes

Opportunity rover used: no
Hours: 69.9
Beginning charge: N/A
Ending charge: N/A
Currently charging: yes

Curiosity rover used: yes
Hours: 138.8
Beginning charge: 100%
Ending charge: 58%
Currently charging: yes

Notes on rovers: N/A

ATV’s Used: yes
Reason for use: Run during the engineer EVA for the battery.
Oil Added? no
ATV Fuel Used: N/A
Hours the ATVs were Used today: 10 minutes
Notes on ATVs: The second one and the red one don’t have anymore battery.

HabCar used and why, where? The habcar is now in the shop awaiting repairs.

CrewCar used and why, where? N/A

General notes and comments: N/A

Summary of internet: All ok.

Summary of suits and radios: Suits 8, 9, 2 and 10 used and radio tested
for the EVA. The ventilation of the suit number 7 wasn’t working. The
light was ON but no air was flowing. Radio number 1 is very noisy, we
can’t hear what the person say.

Summary of Hab operations: N/A

Summary of GreenHab operations: N/A

Summary of ScienceDome operations: N/A

Summary of RAM operations: N/A

Summary of any observatory issues: N/A

Summary of health and safety issues: N/A

Questions, concerns and requests to Mission Support: N/A

Astronomy Report – February 25

Astronomy Report

Name: Eishi Kim
Crew: 222
Date: 2/25/2020

MDRS ROBOTIC OBSERVATORY

Robotic Telescope Requested (choose one) MDRS-WF

Objects to be Imaged this Evening: The double cluster NGC 884/869 and
M45 join the observation queue. The 3 missing Halpha frames of NGC 281
from Sol 4 should also be imaged.

Images submitted with this report: None

Problems Encountered: None

MUSK OBSERVATORY

Solar Features Observed: N/A

Images submitted with this report: N/A

Problems Encountered: N/A

Journalist Report – February 25

Crew 222 Journalist Report 25Feb2020

Author: Marie Bochard, Crew Journalist

Sol 9

Title: More exploration!

This morning, Valentin, Nicolas, Benjamin and I went on EVA to Reservoir Dam up North and Cowboy Corner. After the rain on Saturday night, we wanted to see if there was any damage done to the terrain. We took the rovers Curiosity and Spirit up the path, after re-installing the LOAC experiment close to the Hab. We stopped near the reservoir and discovered that there was barely any water left. This reservoir was man-made to help animals nearby find water, as we deducted from the regularity of the hill. But it hadn’t rained enough to fill it up at all, so we decided to explore the other side of the road. We went down into a small canyon, a tiny replica of the Grand Canyon we visited before coming here. There was a small lake there, with water redder than the sand around it! We found traces of life: there were some antelope footprints on the ground that had dried up after the rain, and leftovers of a cow’s passage (as a reminder, the main road here is call Cow Dung Road… for a reason). My camera’s battery died on me during the EVA, but thankfully we had brought a GoPro to film the landscape, so we got beautiful videos of the exploration.

It felt great being able to leave the Hab again to go exploring. These past few days, we have all been stuck inside because of the mud, only leaving the Hab to change MegaARES’ battery. Tomorrow, it’s Eishi, Nina and Bleuenn’s turn to head out!

For lunch, we had our most anticipated meal: the only pasta we have for the whole mission. We always try to satisfy everyone’s culinary taste, and pasta with pesto was one of the easiest meals we cooked. It brought a taste of home! Now, back to baking bread.

Bleuenn and Benjamin planted new seeds in the GreenHab, and harvested another batch of tiny carrots! The menu keeps improving day after day.

EVA Report – February 25

Crew 222 EVA Report 25-02-2020
EVA #10
Author: Valentin Prudhomme (Engineer)
Purpose of EVA: Replacement of the battery of MegaAres, Setting up the
LOAC and exploration of the reservoir.
Start time: 10:00am
End time: 12:12pm

Narrative: Everything went just fine. We rapidly changed the battery and
assured the waterproofness of the experiment. We installed the LOAC. We
went to the Reservoir Dam. There was very little water. After this we
went to the Cowboy Corner. We found a canyon and went along it.

Destination: Experiments Site and Reservoir Dam.
Coordinates (use UTM NAD27 CONUS): 12S 0518494-4250679 and 12S
0518500-4252800
Participants: Valentin Prudhomme (Engineer, EVA Leader), Benjamin Auzou
(Commander), Marie Bochard (Journalist) and Nicolas Bourdeaud (HSO)
Road(s) and routes per MDRS Map: Entrance road – Cow Dung Road
Mode of travel: Rover – Curiosity and Spirit

GreenHab Report – February 25

Crew 222 GreenHab Report – 25/02/2020

GreenHab Officer : Bleuenn ROIRON

Environmental control : Both heating and ambiant, window open in the
afternoon

Average temperatures: 22,2 °C

Hours of supplemental light: NA

Daily water usage for crops: 13,5 gallons

Daily water usage for research and/or other purposes: 0 Liters

Water in blue Tank: 58,5 gallons

Time(s) of watering crops: 3

Changes to crops: replanted in 3 separated pots white onions, replanted
in 2 separated pots salad, and gave air to raddishes in the big blue pot

Narrative:Today, I started as usual by exposing watercress to UV and
music, and then watered plants. I harvested some basil this morning for
our lunch to add to our pesto. This afternoon I went with Nina and
Benjamin to make some changes to the plants. We also harvested some sun
carrots and planted new carrots, flowers and basil. I also watered
tomatoes again, and came back tonight to water all plants.

Harvest:
Basil: 17g
Sun carrots: 105g(14 little carrots with leaves)

Support/supplies needed: NA

Sol Summary – February 25

Crew 222 Sol Summary Report 25-02-2020
Sol: 9
Summary Title: Liberative day
Author’s name: Benjamin Auzou
Mission Status: all things running

Sol Activity Summary: In the morning, the crew woke up with a matinal
workout session. Then four of us went on EVA to set the LOAC back to its
location. The rainy days are behind us and now all experiments are back
on the field.
In the beginning of the afternoon, some of us took a short nap while the
others were in the solar observatory or in the GreenHab to work on the
plants.
The day was very liberative because we went out for exploration for the
first time in 4 days and also because our HSO prepared us a very
liberative activity in the end of the afternoon in which we danced and
shared a moment to be more confident with ourselves and the others.

Look Ahead Plan:
Tomorrow we are plannning a long EVA. ALso we will start to prepare the
arrival of crew 223, which will continue most of the experiments we are
leading.

Anomalies in work: none
Weather: sunny and windy
Crew Physical Status: all members safe and sound
EVA: LOAC back to experiment site and exploration of the North, between
Reservoir Dam and Cowboy corner.
Reports to be filed: Journalist report, Operations report, GreenHab
Report, Astronomy Report, EVA Report, EVA Request, Photos Report, Photo
of the day
Support Requested: nothing to report